### v1

Abstract
Recent advancements in Artificial Intelligence (AI) demonstrate significant potential to revolutionize weather forecasting. However, most AI-driven models rely on Numerical Weather Prediction (NWP) systems for initial condition preparation, which often consumes hours on supercomputers. Here we introduce XiChen, the first observation-scalable fully AI-driven global weather forecasting system, whose entire pipeline, from Data Assimilation (DA) to medium-range forecasting, can be accomplished within only 17 seconds. XiChen is built upon a foundation model that is pre-trained for weather forecasting. Meanwhile, this model is subsequently fine-tuned to serve as both observation operators and DA models, thereby scalably assimilating conventional and raw satellite observations. Furthermore, the integration of four-dimensional variational knowledge ensures that XiChen’s DA and medium-range forecasting accuracy rivals that of operational NWP systems, amazingly achieving a skillful forecasting lead time exceeding 8.25 days. These findings demonstrate that XiChen holds strong potential toward fully AI-driven weather forecasting independent of NWP systems.
keywords:
fully AI-driven global weather forecasting, deep learning, four-dimensional variational, data assimilation, foundation model
1 Introduction
A century ago, Abbe [1] and Bjerknes [2] conceptualized weather forecasting as an initial value problem, where forecasts could be derived by starting from initial conditions estimated from observations and applying physical laws. At that time, atmospheric observations were extremely limited, computers did not exist, and the predictability of weather remained largely unknown. Thus, their proposition was remarkably ambitious [3]. But today, advancements in our understanding of atmospheric mechanisms, the continuous accumulation of observational data, and significant improvements in computational power have collectively enhanced the accuracy of Numerical Weather Prediction (NWP), making it an indispensable part of daily life [4].
NWP consists of two core components. The first is Data Assimilation (DA), which estimates the most likely state of the atmosphere based on recent observational data from satellites, meteorological stations, ships, and other sources to provide accurate initial conditions for forecasting [3, 4]. The second component is the forecast model, which traditionally simulates the spatial-temporal evolution of atmospheric variables by solving a set of complex Partial Differential Equations (PDEs) [3, 4]. Although NWP has undergone a technological revolution [3], several challenges have emerged. The deceleration of Moore’s Law, increasing complexity of modern NWP systems, and substantial improvements in resolution have made it difficult to deliver timely forecasting results within limited computational time [4, 5]. These limitations consequently impede further progress in NWP.
In recent years, the rapid development of Artificial Intelligence (AI) technology has triggered a transformative revolution in weather forecasting [6]. On dedicated computing hardware, such as Graph Processing Units (GPUs), AI-based models—such as FourCastNet [7], Pangu-Weather [5], GraphCast [8], FengWu [9], Fuxi [10], and AIFS [11]—achieved acceleration rates three to four orders of magnitude faster than operational NWP systems while maintaining comparable forecasting performance. However, these AI models still rely on the (re)analysis fields produced by NWP systems to generate forecasts and are far from a fully AI-driven global weather forecasting system [12, 13, 14]. For instance, the analysis fields used by current AI-based forecasting models (e.g., ECMWF’s AIFS [11]) are products generated through the Four-Dimensional Variational (4DVar) method [15]. The 4DVar method employs optimal control theory to assimilate over 10 million observational data points within each 12-hour Data Assimilation Window (DAW) [16]. Notably, it effectively integrates indirect satellite observations, addressing gaps in atmospheric field information that are challenging to capture using sparse conventional observations [16]. Nevertheless, the 4DVar method in operational NWP systems requires hours of computation on supercomputers [17], restricting the volume of observations that can be assimilated within a limited time frame, with only 5-10% of the available satellite data being used [3]. Additionally, 4DVar requires meticulous tuning and precise formulation of a complex background error covariance matrix [18]. It also depends on sophisticated observation operators [19], tangent linear models, and adjoint models [18], all of which demand extensive development time and expertise from highly trained specialists.
In this era of rapid advancements in AI, a bold proposition has emerged: the development of data-driven systems based on AI technology to forecast future weather states [20, 21] without dependence upon the initial fields from NWP systems. Notably, models such as Aardvark [22] and GraphDOP [23] have successfully generated forecasts directly from raw observations, demonstrating promising performance. However, the natural sparsity of observations poses significant challenges to achieving forecast accuracy comparable to operational NWP systems. Moreover, directly taking all observations as inputs without rigorous quality control may compromise forecast accuracy. Recent studies have also explored integrating pre-trained AI-based weather forecasting models into traditional 4DVar frameworks by leveraging automatic differentiation to create hybrid 4DVar systems [24, 25, 26]. Nevertheless, these efforts often rely on idealized models or simulated observations. Even when real-world conventional observations are employed, these approaches still overlook satellite observations, which are critical for operational NWP systems. On the other hand, some researchers have also ambitiously pursued the development of data-driven DA systems [27, 28, 29, 30, 31], such as FengWu-Adas [32], FuXi-DA [13], DiffDA [30], and 4DVarFormer [14, 33]. These models utilize modern neural network architectures to assimilate conventional and satellite observations, integrating them with AI-based weather forecasting models to create fully AI-driven global weather forecasting systems. Despite these advancements, they can only assimilate observation types encountered during training, necessitating retraining and architectural refinement of the entire model to incorporate new observation types. This limitation restricts their scalability compared to traditional DA systems. Consequently, developing fully AI-driven global weather forecasting systems suitable for operational implementation remains an unresolved challenge.
Here, we propose XiChen (Chinese: {CJK}UTF8gbsn羲辰), the first observation-scalable fully AI-driven global weather forecasting system. XiChen is constructed by a foundation model [34] trained on diverse weather data to replace core components of the NWP pipeline. The foundation model is initially pre-trained on ERA5 [35] data to learn fundamental atmospheric dynamical representations as the medium-range weather forecasting model. Subsequently, it is fine-tuned to incorporate the observation operators and the DA capabilities. By incorporating the 4DVar knowledge into the DA model, specifically by leveraging the gradient of the AI-driven 4DVar cost function as model inputs, we propose a cascaded sequential DA framework to effectively assimilate both conventional and raw satellite observations for initializing accurate global weather forecasting. This cascaded DA framework shows the potential to address the compatibility and robustness challenges associated with the future integration of new observations by fine-tuning. In our experiments, XiChen assimilates conventional observations from the Global Data Assimilation System (GDAS) and raw satellite observational data from the Advanced Microwave Sounding Unit-A (AMSU-A) and the Microwave Humidity Sensor (MHS) within 5 seconds. Although XiChen operates on a global grid with a resolution of 1.40625, its accuracy is comparable to that of operational NWP systems across multiple variables in DA and medium-range weather forecasting tasks. Furthermore, the entire pipeline of XiChen, from DA to 10-day medium-range forecasting, can be accomplished within only 17 seconds using a single A100 GPU, achieving over 400-fold acceleration vs the operational NWP system [17]. Notably, XiChen achieves a skillful forecast lead time of over 8.25 days, which is comparable to the Global Forecasting System (GFS) while surpassing the State-Of-The-Art (SOTA) end-to-end global weather forecasting systems: Aardvark [22] (less than 8 days) and GraphDOP [23] (about 5 days). Furthermore, we propose a dual DA framework to circumvent the multi-hour latency characteristic of the Integrated Forecasting System (IFS HRES), thereby enabling real-time acquisition of medium-range forecast products. We comprehensively evaluate various DA observation configurations and their impact on medium-range forecasting, the effect of 4DVar knowledge on DA and forecasting, and XiChen’s performance in the Tropical Cyclone (TC) track task. These findings highlight that XiChen offers a novel avenue for advancing fully AI-driven global weather forecasting through fine-tuning to scalably integrate new observations in the future.
2 Results
Our experiments evaluated two key aspects of XiChen’s performance: (1) the accuracy and stability of its one-year DA cycle, and (2) the skill of its 10-day medium-range weather forecasts. The DA process employed a cascade framework, sequentially assimilating GDAS prepbufr data, followed by AMSU-A and MHS observations. For the DA cycle experiment, XiChen was initialized at 00:00 UTC on 1 January 2023. The background field was generated by initializing the XiChen forecast model with ERA5 reanalysis data from 00:00 UTC on 27 December 2022, followed by a 5-step rollout with a 24-hour lead time. This forecast field served as the initial condition for the DA cycle, which ran for one year with assimilations at 00:00 and 12:00 UTC daily. We employ a 12-hour DAW wherein observations at 0, 3, 6, and 9 hours after the analysis time are assimilated to initialize the forecasting model for generating the background field for the subsequent DA cycle. This computational process is executed within 5 seconds utilizing a single A100 GPU. Furthermore, we implement a 3-hour DAW that exclusively assimilates observations at the analysis time point to establish the initial conditions for medium-range forecasting. This dual DA framework ensures the fidelity of the assimilation process and for medium-range predictions. Furthermore, XiChen circumvents the multi-hour latency characteristic of IFS, thereby enabling real-time acquisition of medium-range forecast products. Further details of the cascade process and DA cycle are illustrated in Figures S1(a) and S1(b). The weather forecasting model, DA models, and observation operators share the same underlying model architecture and are trained using different data. Details regarding this architecture are provided in Section 4.
For medium-range weather forecasting, we adopted configurations and evaluation metrics from WeatherBench [36] and DABench [33] to generate initial fields and evaluate XiChen’s 10-day forecasts. For details, please see Section 4.8 and Figure S1(c). XiChen’s performance was compared against SOTA operational NWP systems: IFS HRES and GFS. We assessed eight ECMWF Scorecard variables: geopotential at 500 hPa (Z500) and 850 hPa (Z850), 2-meter temperature (T2M), temperature at 850 hPa (T850), Mean Sea Level Pressure (MSLP), 10-meter zonal (10U) and meridional (10V) winds, and specific humidity at 500 hPa (Q500).
Additionally, we evaluated XiChen’s TC assimilation and forecasting capabilities. TC tracks were identified using the TempestExtremes [37] tracker, configured as in NeuralGCM [38], and compared against the International Best Track Archive for Climate Stewardship (IBTrACS) and ERA5. Tracker parameters are detailed in Table S7, and TC track results from forecasts initialized with different observations are shown in Figures S10 to S13. Finally, we present visualizations of the forecasting results produced by XiChen, IFS HRES, and GFS in Figures S14 to S19.
2.1 One year DA Cycle
The time series of the globally averaged, annual latitude-weighted Root Mean Square Error (RMSE) and Bias for the DA cycle results of XiChen, compared to the operational analysis fields from the GFS and the IFS HRES, are presented in Figure 1. Each subplot corresponds to a variable, and skill (y-axis) is plotted at 12-hour intervals across the 365-day horizons (x-axis). Rows 1 and 3 show RMSE, while rows 2 and 4 show Bias. The original data is represented with reduced opacity to enhance clarity, while solid lines depict values smoothed using an Exponential Moving Average (EMA) with a 15-point window. XiChen demonstrates stable operation over long-term DA cycles, with the RMSE of key variables such as Z500, T850, Z850, T2M, U10, V10, and MSLP outperforming those of GFS and approaching those of IFS HRES. Furthermore, the Bias of XiChen remains closer to zero than GFS throughout the year, highlighting its capability to mitigate the bias drift of the forecasting fields. Since the initial background field is a 5-day forecast initialized using ERA5, there is a significant difference compared to the reference. XiChen achieves stable analysis performance after a spin-up period, similar to the 4DVar approach, which can be attributed to integrating 4DVar knowledge within the XiChen DA framework.
The occasional increase in RMSE, as shown in Figure 1, is primarily caused by missing observations. However, it is evident that XiChen effectively mitigates this issue through subsequent assimilation processes. The assimilation is performed using a cascading approach. When AMSU-A or MHS observations are unavailable, the XiChen system excludes the assimilation of AMSU-A or MHS data. Despite this, the accuracy of the analysis fields remains stable. These findings indicate that XiChen can potentially develop into an operational system.
Figure 1: Comparison of the average latitude RMSE (first and third rows) and Bias (second and fourth rows) of the one-year DA cycle using XiChen (red lines) as well as the operational analysis fields of GFS (yellow lines) and IFS HRES (blue lines) for 4 surface variables, such as MSLP, T2M, U10, and V10, and 4 upper-air variables, including Z500, T850, Z850, and Q500, using testing data during 2023. All analysis fields are evaluated against the ERA5 reanalysis dataset.
2.2 Medium-range weather forecasting
The primary metric for evaluating the fully AI-driven global weather forecasting system is its performance in medium-range weather forecasts. In this study, we assess the 6-hour interval medium-range forecasts produced by XiChen. The initial fields for the medium-range forecasting experiment are derived from the analysis fields generated within the DA cycle experiment. The timing of these initial fields is selected based on the setups of WeatherBench [36] and DABench [33]. Figure 2 illustrates the performance of the medium-range forecast initialized by the analysis fields obtained from assimilating GDAS prepbufr observations. Each subplot corresponds to a variable and skill (y-axis) is plotted at 6-hour steps over 10-day horizons (x-axis). Rows 1 and 3 show RMSE, rows 2 and 4 show the Anomaly Correlation Coefficient (ACC). Notably, for key variables such as T2M, T850, MSLP, U10, and V10, XiChen delivers forecasts with lower RMSEs compared to the GFS for nearly the entire 10-day lead time. The RMSEs for all variables approach or even surpass the performance of the IFS HRES as the lead time increases. For Q500, XiChen exceeds the GFS after 2 days and the IFS HRES after 4 days. Although the forecast error for Z500 remains higher than that of both the GFS and the IFS HRES during the first 6 days, XiChen surpasses the GFS after 7 days and outperforms IFS HRES after 8 days.
Figure 2: Comparison of the average latitude RMSE (first and third rows) and ACC (second and fourth rows) of the 10-day medium-range weather forecasting using XiChen (red lines) as well as the operational forecasting results of GFS (yellow lines) and IFS HRES (blue lines) for 4 surface variables, such as MSLP, T2M, U10, and V10, and 4 upper-air variables, including Z500, T850, Z850, and Q500, using testing data during 2023. All analysis fields are evaluated against the ERA5 reanalysis dataset.
Overall, XiChen achieves a skillful weather forecasting lead time of over 8.25 days (with ACC of Z500 $>0.6$), thereby achieving performance comparable to GFS (8.25 days) while substantially outperforming both Aardvark (less than 8 days) [22] and GraphDOP (5 days) [23]. This achievement is particularly noteworthy considering that XiChen was trained using only 12 years of ERA5 data at a 1.40625 resolution. Moreover, XiChen assimilated a limited subset of observations used in operational NWP systems–including GDAS prepbufr, AMSU-A, and MHS observations–significantly fewer than those used by GFS, Aardvark [22], and GraphDOP [23]. XiChen’s RMSE surpasses that of IFS HRES at larger lead times primarily because its forecasts tend to favor smoothing, as evidenced by the energy spectra presented in Figure S8. This tendency is a common characteristic observed in contemporary AI-based weather forecasting models [8, 9, 10]. When initialized using ERA5 data, XiChen demonstrates a skillful forecast lead time of more than 10 days. Further details can be found in the Supplementary Text and Figure S5.
2.3 Contribution of assimilating satellites
As illustrated in Figure S1, XiChen employs a cascade DA framework, which provides the flexibility to select specific observations for assimilation to evaluate the impact of individual satellite observations on the performance of the analysis field. Notably, since the GDAS prepbufr data includes observations for all variables, our approach begins by assimilating GDAS prepbufr observations to effectively estimate the entire atmospheric field. Subsequently, satellite observations are assimilated to further refine the corresponding key variables.
[…]
Figure 3(a) illustrates the impact of separately assimilating AMSU-A or MHS, as well as simultaneously assimilating these two satellite observations, on the analysis quality of upper-air variables, based on the assimilation of GDAS prepbufr data. The first column in each scorecard represents the RMSEs of the analysis field when only GDAS prepbufr data is assimilated. The second column shows the results of additional assimilation of AMSU-A, while the third column reflects the impact of assimilating MHS. The fourth column represents the outcome of assimilating AMSU-A followed by MHS, and the fifth column shows the results of assimilating MHS followed by AMSU-A.
The assimilation of AMSU-A channels 5 to 10, which are primarily designed to observe atmospheric temperature profiles [39], directly improved the upper-air temperature fields. Furthermore, the assimilation of AMSU-A data effectively corrected the upper-air geopotential fields and positively influenced the wind fields. This outcome is likely due to changes in the temperature implicitly adjusting the pressure gradient force, which subsequently modifies the geopotential and wind fields. These findings suggest that improvements in the temperature propagate to the geopotential through their dynamic relationship.
Similarly, the assimilation of channels 3 to 5 of the MHS, intended primarily for observing atmospheric humidity profiles [39], directly enhances the upper-air specific humidity. However, this process adversely affects the upper-air temperature and geopotential fields. This is likely because MHS brightness temperature observations are influenced by factors such as clouds, precipitation, and surface radiation, which are not currently accounted for by XiChen. Furthermore, the extensive adjustments in specific humidity introduced during the DA process may create transient imbalances that disrupt the dynamical state of the atmosphere. Such imbalances often manifest as gravity waves or other types of dynamical feedback, causing the analysis states of the geopotential and wind fields to diverge from equilibrium and thereby amplifying errors. In the future, the incorporation of physical equilibrium constraints into XiChen may be considered as a potential solution to this problem.
Nevertheless, the simultaneous assimilation of AMSU-A and MHS effectively coordinates upper-air atmospheric fields, reducing analysis errors caused by the sparse upper-air observations in GDAS prepbufr. Additionally, the joint assimilation of AMSU-A and MHS mitigates the adverse effects of MHS-only assimilation on the atmospheric temperature and geopotential fields. Importantly, changes in the assimilation order of AMSU-A and MHS have minimal impact on the quality of the analysis fields, highlighting the robustness and scalability of XiChen.
Figure 3(b) presents the medium-range weather forecast scorecard comparing the additional assimilation of two satellite observations, AMSU-A and MHS, to the baseline that uses only GDAS prepbufr observations. The scorecard displays the RMSE values for the corresponding variables at specific pressure levels. In terms of RMSE, the additional assimilation of AMSU-A and MHS demonstrates improved performance across nearly all lead times for upper-air variables, particularly for geopotential, temperature, and specific humidity. These results align with findings from the DA cycle experiment, suggesting that incorporating these satellite observations effectively enhances the analysis of corresponding variables and positively impacts subsequent forecasts. At the pressure level of 50 hPa, forecasts incorporating both AMSU-A and MHS data show significant improvement. This enhancement is largely due to the sparsity of GDAS prepbufr conventional observations in the upper atmosphere. Satellite observations effectively address this data gap, improving medium-range weather forecasts in upper atmospheric levels. For results on the effect of additional assimilation only of AMSU-A or MHS on the weather forecasts after assimilating GDAS prepbufr, see Figure S6 and Figure S7.
2.4 Contribution of the 4DVar knowledge
In this section, we investigate the role of 4DVar knowledge and examine the impact of the trained XiChen model on DA and forecasting performance under different DAWs. In our experiments, the DAW is defined as left-closed and right-open. For instance, a 3-hour DAW incorporates observations only at the initial time, while a 12-hour DAW incorporates observations at the initial time and three additional observations at 3-hour intervals thereafter. We conducted experiments with 3-hour DAW (observations at [0h]), 6-hour DAW (observations at [0h, 3h]), 9-hour DAW(observations at [0h, 3h, 6h]), and 12-hour DAW (observations at [0h, 3h. 6h. 9h]) to evaluate the effect of incorporating 4DVar knowledge over varying periods using the 4DVar cost function gradients.
Figure 4: Comparison of the average latitude RMSE (first and third rows) and Bias (second and fourth rows) of the one-year DA cycle using 3-hourly DAW (blue lines), 6-houly DAW (green lines), 9-hourly DAW (yellow lines), and 12-hourly DAW (red lines) for 4 surface variables, such as MSLP, T2M, U10, and V10, and 4 upper-air variables, including Z500, T850, T500, Z850, and Q500, using testing data during 2023. All analysis fields are evaluated against the ERA5 reanalysis dataset.
We performed a one-year DA cycle experiment, with results presented in Figure 4. The findings indicate that as the DAW length increases, the analysis error decreases, highlighting the positive impact of including more temporal observations on XiChen’s performance. Notably, even without the 4DVar knowledge constraint (DAW [0h]), XiChen maintained a stable DA cycle over one year, despite producing larger analysis errors. This result further demonstrates the robustness of the XiChen model. Additionally, we conducted a 10-day medium-range weather forecast experiment, with results provided in Figure S9. Forecasts generated from analysis fields with longer DAWs exhibited lower RMSE and higher ACC, further confirming that incorporating longer temporal information through 4DVar knowledge enhances forecasting accuracy. These results collectively demonstrate the effective improvement in analysis and forecasting performance achieved by introducing 4DVar knowledge, thereby pushing the fully AI-driven global weather forecasting system towards operationalization.
2.5 Assimilation and forecasting of tropical cyclones
To evaluate the track performance of TCs using XiChen, the IBTrACS is employed as the reference dataset. IBTrACS, accessible at here, consolidates recent and historical TC data from multiple agencies, providing the most comprehensive global collection of TC information.
To assess XiChen’s ability to assimilate and forecast TC trajectories, we employ the TempestExtremes tracker [37] to identify TC tracks. For comparison with the IBTrACS dataset and ERA5 reanalysis, we adopt the configuration used in NeuralGCM [38], with parameter details listed in Table S7. To maintain consistency with IBTrACS, we exclude any TCs identified by this configuration that are not present in the IBTrACS dataset, ensuring that only IBTrACS TCs are included in our results. For details, please see the Supplementary Text.
Figure 5: Assimilation and Forecasting of Tropical Cyclones. (a) Tropical cyclone (TC) trajectory tracks from IBTrACS ground truth data for the entire year of 2023. (b) TC trajectory tracks (blue lines) from the ERA5 reanalysis for 2023. (c) TC trajectory tracks (blue lines) from the XiChen analysis fields obtained by assimilating GDAS prepbufr, AMSU-A, and MHS during a one-year data assimilation (DA) cycle. (d) TC trajectory tracks (blue lines) from the XiChen analysis fields obtained by assimilating GDAS prepbufr and MHS during a one-year DA cycle. (e) TC trajectory tracks (blue lines) from the XiChen analysis fields obtained by assimilating GDAS prepbufr and AMSU-A during a one-year DA cycle. (f) TC trajectory tracks (blue lines) from the XiChen analysis fields obtained by assimilating only GDAS prepbufr during a one-year DA cycle. In panels (b) to (f), the IBTrACS ground truth tracks are shown in gray. (g) Comparison of XiChen performance in terms of mean direct position error over 5-day TC forecasts in 2023. (h) Comparison of XiChen performance in terms of average maximum wind speed error over 5-day TC forecasts in 2023. (i) Comparison of XiChen performance in terms of average minimum MSLP error over 5-day TC forecasts in 2023. In panels (g)-(i), blue lines represent results initialized by assimilating GDAS prepbufr only, while yellow, green, and red lines represent results from additionally assimilating AMSU-A, MHS, and both AMSU-A and MHS, respectively.
This study evaluates XiChen’s performance in assimilating and forecasting the trajectories of TCs. Figure 5(a)-(f) provides a comparative analysis of TC tracks from different datasets and XiChen’s analysis fields. Figure 5(a) depicts the TC trajectories from the IBTrACS ground truth dataset for 2023, including 81 TCs. The blue lines in Figure 5(b) depict the TC trajectories identified in the ERA5 reanalysis dataset for 2023 that correspond with TC trajectories in IBTrack, comprising a total of 52 tropical cyclones. The ERA5 dataset successfully identified 64.2% of the TCs documented in IBTrack, with an average trajectory deviation of 0.66 degrees. The blue lines in Figures 5(c)-(f) represent TC trajectories generated by XiChen during the one-year DA cycle. These trajectories were obtained through various assimilation configurations: GDAS prepbufr with AMSU-A and MHS; GDAS prepbufr with MHS; GDAS prepbufr with AMSU-A; or GDAS prepbufr alone. XiChen identified one more TC following the assimilation of AMSU-A data compared to the assimilation of only GDAS prepbufr data. This improvement is attributed to the enhanced temperature field resulting from AMSU-A assimilation, which subsequently enhances the geopotential field. In turn, this indirectly improves the wind field through a dynamic equilibrium relationship. A more accurate wind field analysis facilitates the identification of closed circulations associated with low-pressure centers, thereby allowing for more precise identification of TCs. Conversely, the number of TCs identified through the additional assimilation of MHS decreased due to its negative impact on the temperature and wind fields. Although the number of identified TCs varies slightly across different observation assimilation scenarios, the overall TC distribution patterns remain consistent, with comparable average trajectory errors. This consistency demonstrates XiChen’s robust performance. Notably, the TC trajectories generated by XiChen include some TCs not identified in ERA5, suggesting that XiChen not only reproduces ERA5 data but also exhibits a degree of generalization capability. Figures S10 through S13 display the results of tropical cyclone trajectory identification from forecasting fields across different forecast lead times.
Furthermore, to assess trajectory and intensity errors in XiChen’s TC forecasts, we initiated 5-day forecasts from the start time of all TCs documented in the IBTrack dataset. The tracking and error calculations were performed at 6-hour intervals. The methodology for tracking TCs in the forecasting fields followed the approach established in the Pangu-Weather [5]. Figures 5(g)–(i) compare four assimilation configurations: (1) assimilating GDAS prepbufr only (blue), (2) assimilating GDAS prepbufr and AMSU‑A (yellow), (3) assimilating GDAS prepbufr and MHS (green), and (4) assimilating GDAS prepbufr, AMSU‑A, and MHS (red). Figure 5(g) shows that assimilating MHS alone increases track errors beyond 48 h, while additional assimilating AMSU‑A mitigates this deterioration. The assimilation of GDAS prepbufr and AMSU‑A does not yield a track improvement relative to GDAS prepbufr only. In Figures 5(h) and 5(i), both maximum wind speed and minimum MSLP errors decrease substantially after assimilating MHS, likely due to improved water‑vapor correction in the atmospheric state. Assimilating AMSU‑A similarly enhances intensity forecasts, and the combined assimilation of AMSU‑A and MHS achieves the greatest reduction in intensity errors. The intensity forecasts after 72 hours exhibit improved accuracy, likely because it is produced by XiChen‑Medium. XiChen‑Medium is fine-tuned using XiChen‑Short’s 5‑day forecast fields as initial inputs, thereby reducing long‑term forecast errors. For details on XiChen‑Short and XiChen‑Medium, see Section 4.4. Notably, our MSLP errors across all lead times closely match those of FuXi-Extreme [40], a medium-range forecasting and post-processing model trained at 0.25 resolution, demonstrating the strong potential of our model for operational use.
3 Disscusion
The development of XiChen represents a significant breakthrough in fully AI-driven global weather forecasting. Leveraging the scalability of foundation models, XiChen is constructed by a pretrained medium-range forecasting model and fine-tuned to serve as observation operators and DA models. The entire pipeline of XiChen, from DA to 10-day weather forecasting, can be accomplished within only 17 seconds, representing an exceeding 400-fold acceleration compared to the operational NWP system [17]. By integrating GDAS prepbufr, AMSU-A, and MHS observations, XiChen offers data-driven global weather forecasting capabilities and demonstrates superior performance compared to the GFS in both DA and forecasting tasks for several key variables. Notably, XiChen achieves a skillful lead time of over 8.25 days, which is comparable to that of the GFS while surpassing Aardvark [22], and GraphDOP [23]. Moreover, XiChen exhibits high scalability: when new satellite observations are introduced, it requires only fine-tuning the foundation model to the respective observation operator, followed by fine-tuning the corresponding assimilation model using the cascade DA framework proposed in this study. Similarly, if an observation becomes temporarily unavailable, the assimilation process can exclude the corresponding model from the pipeline. These capabilities are critical for advancing AI-based DA and forecasting toward practical operationalization. This achievement underscores the transformative potential of AI in advancing operational weather forecasting and highlights the necessity of sustained investments in AI research to tackle the complexities of Earth system modeling. From a multi-agent perspective [41], the XiChen system can be conceptualized as a weak multi-agent system leveraging the pre-training and fine-tuning paradigms of Foundation Models. Within this framework, multiple specialized “agents” are deployed to address distinct tasks: predicting future atmospheric variations based on current states, mapping current atmospheric conditions to satellite observation spaces, and assimilating observations to correct forecast states. Each of these “agents” operates autonomously within its designated function, collectively forming a fully AI-driven global forecasting system that spans the core components of the NWP pipeline—-from assimilating diverse observations to global weather forecasts. As weather forecasting enters a new era, XiChen establishes a robust foundation for applying AI to deepen our forecasting of planetary systems, ultimately guiding decisions with profound implications for society and the environment. However, this does not imply that AI-driven global weather forecasting systems can fully replace NWP. Most AI-driven weather forecasting systems, including XiChen, rely heavily on reanalysis fields produced by NWP systems for training. Consequently, fully AI-driven global weather forecasting, exemplified by XiChen, serves as a valuable complement to NWP rather than a replacement.
Although XiChen demonstrates significant potential for practical applications, it also exhibits certain limitations. First, the model experiences smoothing effects during both the DA and weather forecasting processes, a common challenge in AI-based weather forecasting models [42]. Addressing this issue requires further dedicated research. Potential solutions include designing more suitable loss functions [43], incorporating physical constraints [44], or employing post-processing techniques based on generative models [40]. Second, the current version of XiChen has been trained and tested exclusively at a resolution of 1.40625. To predict finer-scale weather phenomena, future work should focus on extending this resolution to higher levels, such as 0.25 [9] or even 0.09 [45]. Third, the current implementation of XiChen integrates indirect observations from only two satellite instruments, AMSU-A and MHS, which is substantially fewer than the number of satellites used in operational NWP systems [40]. In the future, incorporating additional indirect observational data could significantly enhance the performance of the analysis field. Fourth, when MHS observations are assimilated following the GDAS prepbufr, slight increases in upper-atmosphere temperature and geopotential errors are evident. These discrepancies primarily arise because MHS brightness temperature observations are affected by clouds, precipitation, and surface radiation—factors that XiChen does not currently account for. Additionally, the substantial adjustments to specific humidity during the DA process may induce imbalances. In the future, addressing this issue may involve incorporating additional relevant variables and embedding physical equilibrium constraints into XiChen as potential solutions. Finally, the model produces only deterministic DA and weather forecasts. However, making critical decisions requires awareness of all possible weather scenarios and their associated probabilities. Therefore, it is necessary to extend the model to a probabilistic version. Addressing this limitation in the future may involve reconstructing the model to adopt a probabilistic framework [46] or incorporating perturbations through the deterministic XiChen system [47]. Ultimately, our goal is to develop a fully AI-driven global weather forecasting system that rivals the IFS HRES model in both assimilation and forecasting capabilities. Achieving this ambitious goal necessitates expertise in atmospheric science and the seamless integration of strengths from the atmospheric and AI communities to work together.
4 Medthod
4.1 Notation and problem statement
We denote the full state of the atmosphere at a particular time $t$ as a tensor $\mathbf{x}(t)$ with dimensions $V\times H\times W$, where $V$ represents the total number of variables, $H$ indicates the number of latitude coordinates, and $W$ signifies the number of longitude coordinates. The indexing scheme $\mathbf{x}_{v,i,j}(t)$ refers to the state of variable $v$ at time $t$ and the latitude-longitude coordinates $(i,j)$. Additionally, we represent GDAS prepbufr conventional atmospheric observations at time $t$ as a tensor $\mathbf{y}^{c}(t)$ of the same dimensions $V\times H\times W$. For unobserved variable $v$ at coordinates $(i,j)$, we set $\mathbf{y}^{c}_{v,i,j}(t)=\text{NAN}$, while for observed variable $v$, $\mathbf{y}^{c}_{v,i,j}(t)=\mathbf{x}_{v,i,j}(t)+\mathbf{\varepsilon}_{v,i,j}$, where $\mathbf{\varepsilon}_{v,i,j}$ represents the observation error. For satellite observations, we denote this as $\mathbf{y}^{s}(t)=\mathcal{H}^{s}\left(\mathbf{x}(t)\right)+\mathbf{\varepsilon}^{s}$, where $\mathbf{\varepsilon}^{s}$ is the satellite observation error and $\mathcal{H}^{s}$ is the observation operator that maps atmospheric variables to satellite observation variables.
We aim to develop a fully AI-driven global weather forecasting system that integrates both forecasting and DA models. For the forecasting model, given a system state $\mathbf{x}(t)$, our objective is to predict the state $\mathbf{x}(t+\Delta t)$ at a future time $t+\Delta t$. To accomplish this, the forecasting model $\mathcal{M}_{t\rightarrow t+\Delta t}:\mathbf{x}(t)\mapsto\hat{\mathbf{x}}(t+\Delta t)$ is trained to map the state of the atmosphere at present to a predicted state $\hat{\mathbf{x}}(t+\Delta t)$ for the future time $t+\Delta t$. Regarding the assimilation model, given a background field $\mathbf{x}^{b}(t_{0})$ at time $t_{0}$, GDAS prepbufr observations $\{\mathbf{y}^{c}(t_{0}),\cdots,\mathbf{y}^{c}(t_{K})\}$, and satellite observations $\{\mathbf{y}^{s}(t_{0}),\cdots,\mathbf{y}^{s}(t_{K})\}$ during the DAW $[t_{0},\cdots,t_{K}]$, our goal is to learn to estimate the optimal initial field $\mathbf{x}^{a}(t_{0})$ by fusing these data. To assimilate the satellite observations, we need to learn an observation operator ^$\hat{\mathcal{H}}^{s}:\left(\mathbf{x}(t),\theta^{s}(t)\right)\mapsto\mathbf{y}^{s}(t)$, where $\theta^{s}(t)$ denotes the auxiliary observational information at time $t$. Subsequently, we need to learn a DA model $\mathcal{D}:\left(\mathbf{x}^{b}(t_{0}),\{\mathbf{y}^{c}(t_{0}),\cdots,\mathbf{y}^{c}(t_{K})\},\{\mathbf{y}^{s}(t_{0}),\cdots,\mathbf{y}^{s}(t_{K})\}\right)\mapsto\mathbf{x}^{a}(t_{0})$, where the reanalysis serves as the reference.
4.2 Dataset
In this study, we utilized the DABench [33] dataset while incorporating AMSU-A and MHS satellite observations to train and evaluate the potential of XiChen for satellite assimilation tasks. DABench comprises 14 years of ERA5 reanalysis data, spanning from 2010 to 2023, as well as GDAS prepbufr observations, which were interpolated to a grid resolution of 1.40625 ($128\times 256$ grids). The observations are temporally interpolated to achieve a 3-hour temporal resolution interval, whereby measurements are systematically acquired at 0, 3, 6, 9, 12, 15, 18, and 21 hours UTC daily. Observations obtained within a 1.5-hour temporal window preceding and succeeding these designated temporal reference points are considered temporally coincident. We examined 5 atmospheric variables (each with 9 pressure levels) and four surface variables, resulting in a total of 49 variables. The atmospheric variables include geopotential (Z), temperature (T), specific humidity (Q), the zonal component of wind (U), and the meridional component of wind (V). The 9 sub-variables at different vertical levels are denoted by abbreviating their short names along with their respective pressure levels (e.g., Z500 indicates the geopotential at 500 hPa). The 4 surface variables consist of 2-meter temperature (T2M), 10-meter zonal component of wind (U10), 10-meter meridional component of wind (V10), and mean sea level pressure (MSLP). For further details on the dataset, we refer readers to the article [33].
For the AMSU-A and MHS observations, nearest neighbor interpolation was utilized to map both the raw satellite data and the auxiliary information onto a 1.40625 spatial resolution and 3-hour temporal resolution interval. In this study, we excluded observations from regions located above 60 north and south latitude to avoid complications due to sea ice  [48]. Additionally, we implemented a straightforward quality control procedure that eliminated data with brightness temperature values exceeding 350 K or falling below 150 K.
All models are trained on data collected until 2021, with the year 2022 designated as a validation set. Subsequently, data from 2023 is utilized for testing. For details of the reanalysis dataset as well as observations used in this study, please see Supplementary Text and Tables S1 to S4.
4.3 The fully AI-driven global weather forecasting system
Figure 6: Overview of our XiChen. (a) The overall framework of the fully AI-driven global weather forecasting system “XiChen”. XiChen integrates background field data, conventional observations, and indirect satellite observations within the DAW to perform the DA task. The process begins by computing the 4DVar cost function, and it subsequently determines the gradient with respect to the background field. By combining this gradient with the background field, XiChen generates the analysis field. Utilizing the analysis field along with the lead time as inputs, XiChen produces either a forecast field or an updated background field. The green box denotes the weather forecasting model, the purple represents the observational operator, and the blue signifies the DA model. XiChen undergoes three stages of training: (b) It is pre-trained as a medium-range weather forecasting model, with lead time ($\Delta t$) as the conditional input; (c) It is fine-tuned as an observation operator, using auxiliary observational information of the satellite as the conditional input; (d) It is fine-tuned as a DA model, with the gradient of the 4DVar cost function as the input and the background field as the conditional input; (e) We provide a dual DA framework to operationalize XiChen as a cycling forecasting system.
The XiChen model is based on our proposed Conditional Hybrid Neural Operator (CHNO), which incorporates cross-attention to integrate conditional information into our Hybrid Neural Operator (HNO). This design enables a unified model architecture adaptable to multiple downstream tasks. Additionally, since the Adaptive Fourier Neural Operator (AFNO) [49] performs only global feature extraction through convolution operations in the Fourier space, we introduce a convolution-based branch to enhance local feature extraction capabilities, thereby developing the HNO. For further details on the CHNO and XiChen model architectures, refer to Figure S2. A comparison between HNO and AFNO in the context of medium-range weather forecasting is provided in the Supplementary Text and Figure S4.
Our fully AI-driven global weather forecasting system uses XiChen to cover core components of the NWP pipeline, from observation operators and DA models to the weather forecasting model. As shown in Figure 6(a), XiChen integrates background field data, conventional observations, and satellite observations within the DAW to perform the DA task. The process begins by computing the 4DVar cost function, and it subsequently determines the gradient with respect to the background field. Here, the 4DVar cost function $\mathcal{J}^{\text{4DVar}}(\mathbf{x}(t_{0}))$ represents as follows:
$\tiny{\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))=\frac{1}{2}\|\mathbf{x}^{b}(t_{0})-\mathbf{x}(t_{0})\|^{2}_{\mathbf{B}^{-1}}+\frac{1}{2}\sum^{K}_{k=0}\|\mathbf{y}(t_{k})-\mathcal{H}\left(\mathcal{M}_{t_{0}\rightarrow t_{k}}(\mathbf{x}(t_{0}))\right)\|^{2}_{\mathbf{R}^{-1}},}$ (1)
where $\mathbf{x}(t_{0})$ denotes the initial field to be optimized, $\mathbf{x}^{b}(t_{0})$ represents the background field, and $\mathbf{y}(t_{k})$ represents the observations at time $t_{k}$. The forecasting model, denoted as $\mathcal{M}_{t_{0}\rightarrow t_{k}}$, maps the initial field to the field at time $t_{k}$. Additionally, $\mathbf{B}$ and $\mathbf{R}$ represent the background and observation error covariance matrices, respectively. By combining the gradient $\nabla_{\mathbf{x}^{b}(t_{0})}\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))$ with the background field $\mathbf{x}^{b}(t_{0})$, XiChen generates the analysis field. Utilizing the analysis field along with the lead time as inputs, XiChen produces either a forecast field or an updated background field. Among them, our DA task involves sequentially employing DA models for different observational data. For example, in this study, XiChen first assimilates GDAS prepbufr data, then assimilates AMSU-A observations, and ultimately assimilates MHS observations. This configuration enhances XiChen’s scalability. The introduction of new observations in the future merely requires fine-tuning the respective observation operators and DA model components, eliminating the need to retrain the entire DA system. For details of cascade DA, please refer to the Supplementary Text and Figure S1 (b).
The following subsections will provide a detailed explanation of the medium-range weather forecast pre-training, observation operator fine-tuning, and DA model fine-tuning.
4.4 Pre-training for Medium-range Weather Forecasting
We aim to design a foundation model that can be pre-trained on the reanalysis data and then be finetuned to solve various downstream weather tasks to build a fully AI-driven global weather forecasting system. As illustrated in Figure 6(b), during the pre-training process, the foundation model takes the reanalysis field $\mathbf{x}(t)$ as the input and the lead time $\Delta t$ as the condition and outputs the predicted state $\hat{\mathbf{x}}(t+\Delta t)$. In this study, we sample $\Delta t$ from the set $\{1,3,6,12,24\}$ hours. This process enables the the foundation model to effectively capture the relationships among various weather variables while accounting for both the dynamical and thermodynamic characteristics of the atmosphere. To generate predictions for any time increments, the predictions from the learned forecasting model can be combined using the hierarchical temporal aggregation methodology proposed by Pangu-Weather [5]. Thus, the foundation model is pre-trained to minimize the following loss:
$\tiny{\mathcal{L}(\theta)=\mathbb{E}\left[\frac{1}{VHW}\sum^{V}_{v=1}\sum^{H}_{h=1}\sum^{W}_{w=1}\omega(v)L(i)\|\Delta\mathbf{x}_{v,i,j}(t+\Delta t)\|_{1}\right],}$ (2)
where $\omega(v)$ denotes the pressure-weighted of variable $v$ [8], $\Delta\mathbf{x}(t+\Delta t)=\mathcal{M}_{t\rightarrow t+\Delta t}(\mathbf{x}(t))_{v,i,j}-\mathbf{x}_{v,i,j}(t+\Delta t)$ is the difference between the predicted state and the ground truth at time $t+\Delta t$, $\|\|_{1}$ represents the $L_{1}$-norm, and $L(i)$ is the latitude-weighting factor [36]:
$L(i)=\frac{\cos\left(\text{lat}(i)\right)}{\frac{1}{H}\sum^{H}_{i^{\prime}=1}\cos\left(\text{lat}(i^{\prime})\right)},$ (3)
where $\text{lat}(i)$ is the latitude of the $i$th row of the grid. This term is commonly used in the training process of previous AI-based weather forecasting models [7, 5, 8]. This pre-training process costs about 45 hours to run on 8 A100-40GB GPUs.
To mitigate accumulated errors in the forecasting process, we fine-tune the pre-trained forecasting model by employing a roll-out strategy for multi-step forecasting. Specifically, we execute the model $T$ times for each batch and calculate the average loss of these $T$-step forecasts to optimize the model parameters. The loss function for the multi-step forecast is as follows:
$\tiny{\mathcal{L}(\theta)=\mathbb{E}\left[\frac{1}{TVHW}\sum^{T}_{\tau=1}\sum^{V}_{v=1}\sum^{H}_{h=1}\sum^{W}_{w=1}\omega(v)L(i)\|\Delta\mathbf{x}_{v,i,j}(t+\tau\Delta t)\|_{1}\right],}$ (4)
where $\tau$ denotes the roll out step and $\Delta\mathbf{x}_{v,i,j}(t+\tau\Delta t)=\mathcal{M}_{t\rightarrow t+\tau\Delta t}(\mathbf{x}(t))_{v,i,j}-\mathbf{x}_{v,i,j}(t+\tau\Delta t)$.
In practice, we first fine-tune the pre-trained forecasting model with roll-out steps from $T=2$ to $T=4$ to develop the XiChen-Short model. The weights obtained from the XiChen-Short model are then used to initialize the XiChen-Medium model, which is subsequently fine-tuned to achieve optimal forecasting performance for periods of 5 to 10 roll-out steps. We use the same sampled value of $\Delta t$ for all $T$ steps. Our forecasting model is similar to FuXi [10] in its forecasting process. Since the analysis field errors provided by the XiChen DA model are comparable to those of XiChen-Short’s 2-day forecasts initialized with ERA5, the 10-day medium-range weather forecasts using initial fields estimated by the XiChen DA model are generated by XiChen-Short for the first 3 days and by XiChen-Medium for the remaining 7 days. The fine-tuning process of the XiChen forecasting model costs about 94 hours to run on 8 A100-40GB GPUs. For additional information on the model training configuration, please refer to the Supplementary Text.
4.5 Fine-tuning for the Observation Operator
To compare the atmospheric state with satellite observations, the observation operator is essential [19]. The simulated radiance is computed at each observation location using a profile of atmospheric variables. We fine-tune the pre-trained model by taking the satellite’s attitude parameters and scan position as conditions to construct an observation operator ^$\hat{\mathcal{H}}^{s}:\left(\mathbf{x}(t),\theta^{s}(t)\right)\mapsto\mathbf{y}^{s}(t)$, where $\theta^{s}(t)$ denotes the auxiliary observational information at time $t$. The loss function used to train ^$\hat{\mathcal{H}}^{s}$ is as follows:
$\mathcal{L}=\mathbb{E}\left[\frac{1}{C}\sum^{C}_{c=1}\|\hat{\mathcal{H}}^{s}\left(\mathbf{x}(t),\theta^{s}(t)\right)_{c}-\mathbf{y}^{s}_{c}(t)\|_{1}\right],$ (5)
where $C$ denotes the channels of the satellite observation, and the subscript $c$ represents the channel to be calculated. This process facilitates the development of scalable plug-ins to assimilate the satellite’s raw observations.
This study selected channels 5 to 10 of AMSU-A and 3 to 5 of MHS for training observation operators and conducting assimilation experiments. This selection was motivated by the fact that channels 5 to 10 of AMSU-A are sensitive to atmospheric temperature profiles ranging from 50 hPa to 1000 hPa, while channels 3 to 5 of MHS are responsive to the atmospheric water vapour in the troposphere [39]. The fine-tuning process of each observation operator costs about 13.3 hours to run on 8 A100-40GB GPUs. Please refer to the Supplementary Text for details regarding the configuration of the AMSU-A and MHS observation operators and their performance in Tables S5 and S6.
4.6 Fine-tuning for Data Assimilation
Notably, we leverage the atmospheric dynamics information provided by the XiChen forecast model by incorporating 4D-Var a priori knowledge. This approach effectively propagates observations across the atmospheric field’s spatial and variable dimensions. Furthermore, the 4D-Var cost function enables the seamless integration of diverse observational operators, facilitating the assimilation of raw satellite observations.
As illustrated in Figure 6(c), we fine-tune the pre-trained model to function as a DA model by incorporating the gradient of the 4DVar cost function with respect to the background field, denoted as $\nabla_{\mathbf{x}^{b}(t_{0})}\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))$. This approach follows the derivation presented in 4DVarFormer [14]. However, unlike 4DVarFormer, XiChen outputs the representation of analysis increments in latent space rather than the analysis increments themselves. This representation is added to the embedding of the background field $\mathbf{x}^{b}(t_{0})$ and subsequently decoded by the HNO blocks to generate the analysis field.
In our 4DVar cost function, the observation covariance matrix of prepbufr observations is denoted as $\mathbf{R}^{c}=\varepsilon^{c}(\varepsilon^{c})^{T}$. The observation operator error $\hat{\mathbf{\varepsilon}}^{s}$ is used to construct the error covariance matrix, $\mathbf{R}^{s}=\hat{\varepsilon}^{s}(\hat{\varepsilon}^{s})^{T}$, of the satellite observations. Since we only need to compute the gradient of the initial value of the 4DVar cost function with respect to $\mathbf{x}^{b}$, there is no need to estimate $\mathbf{B}$. This simplifies the process and circumvents a challenging issue typically encountered in the traditional 4DVar method [50]. In addition, unlike methods used in Aardvark [22], GraphDOP [23], FuXi-DA [13], and FengWu-Adas [32], we incorporate a quality control process into the computation of the 4DVar cost function by excluding observations whose distances from the background field exceed 5 times the standard deviation of the variable. This approach can prevent excessive adjustment of localized regions in the background field caused by large discrepancies between observations and the background. For a detailed depiction of the computational flow of the 4DVar cost function, please refer to Figure S3.
Therefore, we fine-tune the pre-trained model by taking $\nabla_{\mathbf{x}^{b}(t_{0})}\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))$ as the input and the background field $\mathbf{x}^{b}$ as the condition to construct an DA model $\mathcal{D}$. The model is trained to minimize the following loss function:
$\mathcal{L}(\theta)=\mathbb{E}\left[\frac{1}{VHW}\sum^{V}_{v=1}\sum^{H}_{i=1}\sum^{W}_{j=1}\omega(v)L(i)\|\left(\mathbf{x}^{a}_{v,i,j}(t)-\mathbf{x}_{v,i,j}(t)\right)\|_{1}\right].$ (6)
We initially fine-tuned the pre-trained model using GDAS prepbufr observations to develop a prepbufr DA model, which costs about 45 hours to run on 8 A100-40GB GPUs. Subsequently, the prepbufr DA model was further fine-tuned into an AMSU-A DA model, employing the analysis field generated by assimilating GDAS prepbufr as the background field, which costs about 94 hours to run on 8 A100-40GB GPUs. Finally, the AMSU-A DA model was sequentially fine-tuned to assimilate MHS satellite observations, with the assimilating order of AMSU-A and MHS randomized, which costs about 222 hours to run on 8 A100-40GB GPUs. Detailed configurations for the assimilation model training are provided in the Supplementary Text.
4.7 Dual DA framework to operationalize XiChen as a cycling forecasting system
As shown in Figure 6(e), to circumvent the temporal latency inherent in AI-based 4DVar methods [24, 14] when deploying XiChen as a continuous forecasting system for acquiring medium-range prognostic results, we have architected a dual assimilation framework. Within this framework, we establish a configuration encompassing a 12-hour DAW and a 3-hour DAW. In the 12-hour DAW, we systematically assimilate observations retrospectively from the analysis time point at intervals of 0, 3, 6, and 9 hours to derive an accurate analysis field. This analysis field serves as the input for XiChen to generate a 12-hour forecast, subsequently yielding the background field for the initialization of the subsequent DAW cycle. Concurrently, within the 3-hour DAW configuration, we assimilate observations contemporaneous with the analysis time to generate the initial condition to initialize XiChen for the provision of medium-range forecast products. This dual DA framework facilitates the real-time acquisition of medium-range weather forecasting results without necessitating the temporal delay associated with the availability of all observations within the complete 12-hour DAW. In comparison to the IFS HRES, XiChen can deliver medium-range forecast products with a temporal advantage of several hours, thereby providing accurate weather forecasting products essential for governmental decision-making processes and people’s travel plans expeditiously.
4.8 Verification metrics
This study aims to assess our model’s performance concerning the tasks of assimilation and forecasting. Consequently, this study comprehensively evaluates a one-year DA cycle and a 10-day medium-range forecast. The DA cycle is executed at 12-hour intervals for assimilations, with a DAW of 12 hours. Specifically, the DA cycle is run for the year 2023 at 00:00 UTC and 12:00 UTC each day, which corresponds to the initialization times for the 10-day forecast conducted by the IFS HRES, Pangu-Weather [5], GraphCast [8], and FengWu [9]. To evaluate medium-range forecasts, we follow the WeatherBench [36] and DABench [33], selecting 50 initial fields at 336-hour intervals for the medium-range forecast experiments. The first initial field at 00:00 UTC is set for January 1, 2023, while the first initial field at 12:00 UTC is on January 8, 2023.
All metrics were computed using float32 precision and reported using the native scale of the variables without normalization. Notably, all metrics are computed using a latitude-weighting factor over grid points due to the non-equal area distribution from the equator towards the north and south poles.
Root Mean Square Error (RMSE)
We evaluate assimilate and forecast skill for a given variable, $\mathbf{x}_{v}$, using a latitude-weighted Root Mean Square Error (RMSE) [36] given by
$\text{RMSE}_{v}=\frac{1}{|D_{eval}|}\sum_{d\in D_{eval}}\sqrt{\frac{1}{HW}\sum^{H}_{i=1}\sum^{W}_{j=1}L(i)(\hat{\mathbf{x}}_{v,i,j}-\mathbf{x}^{t}_{v,i,j})^{2}},$ (7)
where
  • $\hat{\mathbf{x}}$ is the field to be evaluated,
  • $\mathbf{x}^{t}$ is the ERA5 ground truth,
  • $d\in D_{eval}$ represent the sample index in the evaluation dataset,
  • $i$ represents the latitude coordinate in the grid,
  • $j$ represents the longitude coordinate in the grid.
The lower the RMSE, the better the results.
Bias
We also computed the bias for a given variable, $\mathbf{x}_{v}$
$\text{Bias}_{v}=\frac{1}{HW}\sum^{H}_{j}\sum^{W}_{k}L(i)(\hat{\mathbf{x}}_{v,i,j}-\mathbf{x}^{t}_{v,i,j}).$ (8)
The closer the Bias is to 0, the better the results are.
Anomaly Correlation Coefficient (ACC)
To study skillful forecast lead times, we also calculated the latitude-weighted Anomaly Correlation Coefficient (ACC) [36] according to
$\text{ACC}_{v}=\frac{\sum_{d\in D_{eval}}\frac{\sum^{H}_{i}\sum^{W}_{j}L(i)\left(\hat{\mathbf{x}}_{v,i,j}-\mathbf{C}_{v,i,j}\right)\left(\mathbf{x}^{t}_{v,i,j}-\mathbf{C}_{v,i,j}\right)}{\sqrt{\left[\sum^{H}_{i}\sum^{W}_{j}L(i)\left(\hat{\mathbf{x}}_{i,j}-\mathbf{C}_{v,i,j}\right)^{2}\right]\left[\sum^{H}_{i}\sum^{W}_{j}L(i)\left(\mathbf{x}^{t}_{v,i,j}-\mathbf{C}_{v,i,j}\right)^{2}\right]}}}{|D_{eval}|},$ (9)
where $\mathbf{C}_{v,I,j}$ denotes the climatological mean for a given variable $v$ and the day-of-year containing the validity time at the ${i,j}$ grid point. It is calculated referring to GraphCast [8], FengWu [9], and DABench [33]. The hourly climatological fields are computed using ERA5 data between 2010 and 2021. The higher the ACC, the better the results.
[…]
Supplementary information The supplementary material is available at the “Supplementary_Information.pdf”.
[…]
Acknowledgements The authors extend their gratitude to the ECMWF and NCEP for their significant efforts to store and provide invaluable data, which are crucial for this work and the research community. Additionally, this work was carried out at the National Supercomputer Center in Tianjin, and the calculations were performed on Tianhe new generation supercomputer. We would also like to express our appreciation to the research team and service team in the Shanghai Artificial Intelligence Laboratory for the provision of computational resources and infrastructure. This work was supported by the Science and Technology Innovation Program of Hunan Province (2022RC3070) and the National Natural Science Foundation of China (Grant Nos.  42205161, 42405146, and 42275170).
[…]
Funding Not applicable.
[…]
Conflict of interest All authors declare no financial or non-financial competing interests.
[…]
Ethics approval and consent to participate Not applicable.
[…]
Consent for publication Not applicable.
[…]
Availability of data and materials All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Information. The DABench dataset is available at the Baidu Drive. AMSU-A and MHS observations can be accessed here.
[…]
Code availability The source code used for this work is available in a GitHub repository https://github.com/wuxinwang1997/XiChen_1.40625deg.
[…]
Author contribution K.J.R., L.B., B.H.D., and W.X.W. designed the project. K.J.R., L.B., W.C.N., and B.H.D. managed and oversaw the project. W.X.W., S.M., and Y.L.Z. selected and preprocessed the observations. W.X.W., T.K.Y., and T.H. performed the model training and evaluation. W.X.W., T.H., and L.L.H. improved the model design. W.X.W., T.K.Y., X.Y.L., K.F.D., and T.H. analyze the experiment results. W.X.W., B.F., and W.C.N. wrote and revised the manuscript.